OpenStudy (anonymous):

If $240 is invested at an interest rate of 9% per year and is compounded monthly, how much will the investment be worth in 14 years?

OpenStudy (anonymous):

This is a periodic compounding exercise: a yearly interest rate with several compounding within the year. FV = \[FV = PV (1+r/n)^c*n\] r = 9% = 0.09 n = 12 c = 14 0.09 / 12 = 0.0075 Therefore, FV = 240 ( 1 + 0.075) ^ 168 FV = 842.1325